# FlowDreamer: Exploring High Fidelity Text-to-3D Generation Via Rectified Flow

### v2

@@ 4.3 Quantitative Comparisons
We use CLIP (Radford et al., 2021) similarity to quantitatively evaluate our method under either NeRF (Mildenhall et al., 2021) or 3D GS (Kerbl et al., 2023) settings. The results of 3D representation with NeRF come from implementation in  (Guo et al., 2023). The results of 3D representation with 3D GS are from their the official implementation. The prompts of NeRF results are all from Dreamfusion, and the prompts of 3D GS are from LucidDreamer (Liang et al., 2023) and ChatGPT.
@@
We randomly choose 26 prompts each to compare in 3D GS and NeRF. We randomly select 12 images from the rendered images. The rendered images are from azimuth angles from -180 to 180 degrees with a fixed elevation of 15 degrees for both NeRF and 3D GS. We use three CLIP models from OpenCLIP (Ilharco et al., 2021), ViT-B-32, ViT-L-14, and ViT-g-14, to calculate the CLIP similarity. Our method demonstrates better CLIP similarities both in NeRF and in 3D GS scenarios. As shown in Tab. 1, our CLIP similarity achieved the best results across all three CLIP models, with the largest margin of 2.57 over the second-best result in ViT-g-14 in NeRf results. And Tab. 2 shows that our CLIP similarity also achieved the best results compared to other methods. In particular, it exceeds the LucidDreamer result by 1.89 in ViT-B-32.
@@ 4.4 Experimental Insights of our FlowDreamer
Figure 8: Compared to other baseline methods with the same rectified flow prior, our FlowDreamer generates 3D results with richer details and more realistic colors.
@@
3D generation with rectified flow prior. To better demonstrate the effectiveness of our method, we replace the best SOTA method Luciddreamer’s diffusion prior with the Rectified flow prior. The derivation process is provided in the supplementary material. We refer to the ISM loss of LucidDreamer with Rectified Flow Prior as VF-ISM. Figure 8 shows that our method can generate results with finer details and more realistic textures compared with VfDS and VF-ISM. For example, the tank’s tracks and the duck’s feathers are heavily blurred in VFDS. The tank’s color is less realistic, and the color of the duck’s back is somewhat oversaturated in VF-ISM. However, our method not only generates the tank and the duck with richer details but also presents a more realistic appearance.
@@
Faster convergence phenomenon. Our FlowDreamer can be trained to converge faster than our initial framework VFDS. When reaching 1200 steps, our FlowDreamer can generate fine-grained objects whereas VFDS requires more steps to achieve convergence (see Figure 9). When the training step reaches 3000, VFDS still shows much less detail than the FlowDreamer.
Figure 10: A comparison between our VFDS (upper) and FlowDreamer (bottom) with different scales of CFG. Our method shows better generation performance. Prompt: “A cake filled with Oreos, highly detailed, photorealistic.”
Figure 11: Impacts of different NFEs. “NFE N+1” denotes using N steps of iteration for push-backward and 1 step for gradient calculation.
Figure 12: Different sampling methods, where ‘NFE: 2*3+1’ indicates that the second-order method requires 2 inferences per iteration, the total push-backward process takes 3 iterations, and an additional evaluation is needed for gradient calculation.
Impacts of different Number of Function Evaluations (NFE). As NFE increases, our FlowDreamer generates objects with more details and more complex structures (see Figure 11). For example, the structure complexity of the front hood of the LEGO car gradually increases. However, Even with a small NFE, wherein the push-backward process has a small cost, our method can still train a 3D model with good performance.
Impacts of various sample methods. We test three sample methods, namely first-order Eulder, second-order midpoint, and fourth-order Runge-Kutta. Experimental results in Figure 12 show that higher-order solvers do not necessarily yield better performance. For more experimental results, please refer to the supplementary materials. For example, the total push-backward process takes three iterations, and the Euler method actually produces more realistic results.
5 Conclusion
In this paper, we explored a new direction by using the rectified flow model as an alternative prior for text-to-3D generation. We developed a mathematical analysis to adapt SDS to rectified flow model, resulting in the initial VFDS framework. However, VFDS still leads to over-smoothing. We analyzed this issue from the perspective of ODE trajectories and proposed FlowDreamer, a text-to-3D framework with a new UCM loss. Extensive experiments showed that FlowDreamer achieves high-fidelity results with richer details and faster convergence in both NeRF and 3D GS settings. We also highlighted open questions, such as initialization issues for NeRF and noise search sampling.
6 Limitation
The Jabus problem still exists; simply adding words like ‘front view,’ ‘back view,’ and ‘side view’ to the prompt is insufficient for supervising the generation view. Although we attempt to mitigate this issue by using Perpneg (Armandpour et al., 2023) in our FlowDreamer method, the problem still occurs occasionally. We consider solving the Jabus problem thoroughly as a focus for future work.

### v3

@@ 4.2 Quantitative Comparisons
We use CLIP (Radford et al., 2021) similarity to quantitatively evaluate our method under either NeRF (Mildenhall et al., 2021) or 3D GS (Kerbl et al., 2023) settings.
@@
The results of 3D representations with NeRF come from implementation in (Guo et al., 2023). The results of 3D representation with 3D GS are from their official implementation. The prompts of NeRF results are from DreamFusion, and the prompts of 3D GS are from LucidDreamer Liang et al. (2023) and ChatGPT.
We randomly choose 26 prompts each to compare in 3D GS and NeRF. We randomly select 12 from the rendered images. The rendered images are from azimuth angles from -180 to 180 degrees with a fixed elevation of 15 degrees for both NeRF and 3D GS. We use three CLIP models from OpenCLIP (Ilharco et al., 2021), ViT-B-32, ViT-L-14, and ViT-g-14, to calculate the CLIP similarity. Our method demonstrates better CLIP similarities both in NeRF and in 3D GS scenarios.
@@
As shown in Tab. 1, our CLIP similarity achieved the best results across all three CLIP models, with the largest margin of 2.57 over the second-best result in ViT-g-14 in NeRF results. And Tab. 2 shows that our CLIP similarity also achieved the best results compared to other methods. In particular, it exceeds the LucidDreamer result by 1.89 in ViT-B-32.
@@ 4.3 Experimental Insights of our FlowDreamer
Figure 7: Comparison with other baseline methods using the same rectified flow prior under 3D GS and NeRF generation settings. Our 3D results are with richer details and more realistic colors.
@@
3D generation with rectified flow prior. To better demonstrate the effectiveness of our method, we replace the SOTA method LucidDreamer’s diffusion prior with the rectified flow prior. The derivation process is provided in the supplementary material. We refer to the ISM loss of LucidDreamer with vector field as VF-ISM. Figure 7 demonstrates that our method can generate results with finer details and more realistic textures compared with VFDS and VF-ISM in both 3D GS and NeRF results. FlowDreamer achieves convergence in 3D GS and NeRF faster than VFDS or VF-ISM, while demonstrating superior details and more realistic shapes. For instance, Figure 7(a) in 3D GS indicate that the tank’s tracks and the duck’s feathers appear heavily blurred in VFDS. Furthermore, the tank’s color lacks realism, and the duck’s back is somewhat oversaturated in VF-ISM. Our FlowDreamer not only generates the tank and the duck with richer details but also achieves a more realistic overall appearance. In addition, the Figure 7(b) NeRF results reveal that the axe shape and details are improved, whereas the corgi in VFDS exhibits excessive smoothing, and the corgi in VF-ISM presents some noise. The shape and details of our corgi remain relatively satisfactory.
Faster convergence. Our FlowDreamer can be trained to converge faster than the initial framework VFDS. When reaching 1200 steps, our FlowDreamer can generate fine-grained objects whereas VFDS requires more steps to achieve convergence (please refer to the supplementary material). When the training step reaches 3000, VFDS still shows much fewer details than FlowDreamer, and it takes less time overall.
Figure 8: A comparison between our initial framework VFDS (upper) and FlowDreamer (bottom) with different scales of CFG. The results generated by FlowDreamer contain more detailed features. Prompt: “A cake filled with Oreos, highly detailed, photorealistic.”
Figure 9: Impacts of different NFEs and sampling methods. “NFE N+1” denotes using $N$ steps of iteration for push-backward, and 1 step for gradient calculation. ‘NFE: 2*3+1’ indicates that the second-order method requires 2 inferences per iteration. The total push-backward takes 3 iterations.
Impacts of different Number of Function Evaluations (NFE). As NFE increases, the training time also increases, and the generated objects exhibit more details and more complex structures (see Figure 9(a)). For example, the structure complexity of the front hood of the LEGO car gradually increases. However, Even with a small NFE, wherein the push-backward process has a small cost, our method can still train a 3D model with good performance.
Impacts of various sampling methods. We test three sampling methods, namely first-order Eulder, second-order midpoint, and fourth-order Runge-Kutta. Experimental results in Figure 9(b) show that higher-order solvers do not necessarily yield better performance. For example, the total push-backward process takes three iterations, and the Euler method actually produces more realistic results while requiring the least amount of time. For more experimental results, please refer to the supplementary material.
5 Conclusion
In this paper, we explored a new direction by using the rectified flow model as an alternative prior to text-to-3D generation. We developed a mathematical analysis to adapt SDS to rectified flow model, resulting in the initial VFDS framework. However, VFDS still leads to over-smoothing. We analyzed this issue from the perspective of ODE trajectories and proposed FlowDreamer, a text-to-3D framework with a new UCM loss. Extensive experiments showed that FlowDreamer achieves high-fidelity results with richer details and faster convergence in both NeRF and 3D GS settings. We also highlighted open questions, such as initialization issues for NeRF and noise search sampling. Limitation. The Jabus problem still exists; simply adding words like ‘front view,’ ‘back view,’ and ‘side view’ to the prompt is insufficient for supervising the generation view. Although we attempt to mitigate this issue using Perpneg (Armandpour et al., 2023), it still occasionally occurs. We consider solving the Jabus problem thoroughly as a focus for the future work.
Appendix
The overview of the Appendix: We provide some illustrations, including the effects of the Transformer Jacobian term, a comparison of the speeds of SDS and VFDS, a comparison of the speeds of VFDS and FlowDreamer, and an example of the initialization difference between NeRF and 3D GS models (see .1). Accordingly, we introduce the details of the experiments in .2. To better compare with other methods, we also conduct a user study to evaluate user preferences in .3. The derivation process of VF-ISM with the Rectified Flow prior is discussed in .4. Additionally, we discuss the steps of NFE and sampling methods in .5. More results include comparisons under 3D GS and NeRF generation settings and additional results from our FlowDreamer (see .6).
.1 Some illustrations
We provide some illustrations for a better understanding of our paper. Figure 10 illustrates the effects of the Transformer Jacobian term. It is shown that keeping this term leads to training crashes, making it very difficult to generate meaningful 3D objects. Therefore, we omit this term to achieve an effective gradient for optimization. Figure 11 shows a comparison of the convergence speeds of SDS and VFDS. Figure 12 shows a comparison of the convergence speeds of VFDS and FlowDreamer. Figure 13 illustrates an example of the initialization difference between NeRF and 3D GS models.
Figure 10: VFDS results of ignoring the Transformer Jacobian. We found that it is difficult to generate meaningful results.
Figure 11: A comparison of SDS and VFDS with different steps shows that VFDS converges more quickly and achieves relatively good results in just 3000 steps. Prompt: “A lemon cake.”
@@ 5 Conclusion
(a) NeRF
(b) 3D Gaussian splatting
.2 Implementation Details
We adopt Stable Diffusion 3 (SDv3) (Esser et al., 2024) as our Rectified flow model. To facilitate a better comparison, we will categorize the methods into two types: one where the 3D model is NeRF, and the other where the 3D model is 3D GS. For NeRF results comparisons, unlike ProlificDreamer (Wang et al., 2024a) and Consistent3D (Wu et al., 2024), which use multi-stage rendering with normal, geometric, and texture rendering, our method simply uses Instant-NGP (Müller et al., 2022) for rendering. It is optimized with a resolution of 256 for the first 5000 steps with a batch size of 1, and then 512 for the subsequent 3000 steps, also with a batch size of 1. For the warm-up strategy, we use VFDS to optimize over 1200 steps. The initial framework VFDS need more steps to optimize, which uses 5000 steps for both 256 and 512 resolutions, with the batch size always set to 1. For 3D Gaussian splatting comparisons, we utilize the pretrained PointE (Nichol et al., 2022) to initialize the locations of 3D Gaussians, while other properties of 3D Gaussians adopt random initialization. Our FlowDreamer uses a batch size of 4 and 3000 iterations, while VFDS uses the same batch size but with 4000 iterations. The CFG for VFDS is set to 100 on both NeRF and 3D GS, while it is set to 40 for FlowDreamer.
.3 User study
Table 3: User study of 3D Gaussian splatting methods.
Methods DreamGaussian (Tang et al., 2023) GaussianDreamer (Yi et al., 2023) LucidDreamer (Liang et al., 2023) Ours
Scores 1.18 2.03 3.09 3.70
Table 4: User study of NeRF methods.
Methods DreamFusion (Poole et al., 2022) ProlificDreamer (Wang et al., 2024a) Consistent3D (Wu et al., 2024) Ours
Scores 1.46 2.70 2.18 3.65
To compare our method with other comparison methods based on human perception, we conducted a user study involving 36 participants using generated videos from 26 prompts in NeRF and 3D Gaussian splatting, respectively. Participants viewed four videos simultaneously and assigned scores of 1, 2, 3, and 4 to each video. In each test, the prompts were shown in the title, and participants were instructed to make their decisions based on the degree of alignment of the video with the text, the detail of the video, the color of the video, and the quality of the video. A higher score indicates that users believe the video is better. Our FlowDreamer achieves the highest scores among these methods, both in NeRF results and in 3D Gaussian splatting results.
.4 VF-ISM Derivation
ISM uses DDIM inversion to predict the noisy latent $x_{s}$ as below.
$x_{s}$ $=\sqrt{\bar{\alpha}_{s}}\hat{x}^{s-\delta_{T}}+\sqrt{1-\bar{\alpha}_{s}}\epsilon_{\phi}(x_{s-\delta_{T}},s-\delta_{T},\emptyset)$ (9)
$=\sqrt{\bar{\alpha}_{s}}\left(\hat{x}^{s-\delta_{T}}+\gamma(s)\epsilon_{\phi}(x_{s},s,\emptyset)\right)$
where $x^{s-\delta_{T}}=\frac{1}{\sqrt{\bar{\alpha}_{s}}}x_{s-\delta_{T}}-\gamma(s-\delta_{T})\epsilon_{\phi}(x_{s-\delta_{T}},s-\delta_{T},\emptyset)$, $\gamma(t)=\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}$ and $s=t-\delta_{T}$. The $\hat{x}^{s-\delta_{T}}$ is computed using DDIM inference, while $x$ represents the rendered image from the 3D model, as shown below.
$\hat{x}^{s-\delta_{T}}$ $=x-\gamma(\delta_{T})\left[\epsilon_{\phi}(x_{\delta_{T}},\delta_{T},\emptyset)-\epsilon_{\phi}(x,0,\emptyset)\right]\ldots$ (10)
$\quad-\gamma(s-\delta_{T})\left[\epsilon_{\phi}(x_{s-\delta_{T}},s-\delta_{T},\emptyset)-\epsilon_{\phi}(x_{s-2\delta_{T}},s-2\delta_{T},\emptyset)\right]$
Next, ISM computes $x_{t}$ based on $x_{s}$.
$x_{t}$ $=\sqrt{\bar{\alpha}_{t}}\hat{x}^{s}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\phi}(x_{s},s,\emptyset)$ (11)
$\hat{x}^{s}$ $=x^{s-\delta_{T}}-\gamma(s)\left[\epsilon_{\phi}(x_{s},s,\emptyset)-\epsilon_{\phi}(x_{s-\delta_{T}},s-\delta_{T},\emptyset)\right]$
After that, ISM computes the $\tilde{x}^{t}$ with the denoising process.
$\tilde{x}^{t}=\frac{x_{t}}{\sqrt{\bar{\alpha}_{t}}}-$ $\gamma(t)\epsilon_{\phi}(x_{t},t,y)+\gamma(s)\left[\epsilon_{\phi}(x_{t},t,y)-\epsilon_{\phi}(x_{s},s,y)\right]$ (12)
$+...+\gamma(\delta_{T})\left[\epsilon_{\phi}(\tilde{x}_{2\delta_{T}},2\delta_{T},y)-\epsilon_{\phi}(\tilde{x}_{\delta_{T}},\delta_{T},y)\right]$
ISM computes the $x-\tilde{x}^{t}$ and with the DDIM inversion process.
$x-\tilde{x}^{t}$ $=\gamma(t)\left[\epsilon_{\phi}(\tilde{x}_{t},t,y)-\epsilon_{\phi}(x_{s},s,\emptyset)\right]+\eta_{t}$ (13)
[…]
$+\gamma(\delta_{T})\left[\epsilon_{\phi}(\tilde{x}_{\delta_{T}},\delta_{T},y)-\epsilon_{\phi}(x,0,\emptyset)\right]-\gamma(\delta_{T})\left[\epsilon_{\phi}(\tilde{x}_{2\delta_{T}},2\delta_{T},y)-\epsilon_{\phi}(x_{\delta_{T}},\delta_{T},\emptyset)\right]$
To derive VF-ISM, we adapt ISM to rectified flow. Concretlty, we first conduct push-backward operation to compute the $x_{s}$ with Euler sample method as below.
$x_{\delta_{T}}$ $=x+v_{\phi}(x,0,\emptyset)dt$ (14)
$x_{2\delta_{T}}$ $=x_{\delta_{T}}+v_{\phi}(x_{\delta_{T}},\delta_{T},\emptyset)dt$
$x_{s}$ $=x_{s-\delta_{T}}+v_{\phi}(x_{s-\delta_{T}},s-\delta_{T},\emptyset)dt$
where, $dt=\Delta_{T}$.
Then we use the push-backward to compute the $x_{t}$ with Euler method:
$x_{t}=x_{s}+v_{\phi}(x_{s},s,\emptyset)ds$ (15)
where, $ds=t-s$.
Due to rectified flow, the $\tilde{x}_{t}^{t}$ can be simplly expressed as below:
$\tilde{x}^{t}$ $=x_{t}-v_{\phi}(x_{t},t,y)ds$ (16)
$-v_{\phi}(x_{s},s,y)dt-v_{\phi}(x_{s-\delta_{T}},s-\delta_{T},y)dt$
$-...-v_{\phi}(x_{\delta_{T}},\delta_{T},y)dt$
Then we compute the $x-\tilde{x}^{t}$ with push-backward with rectified flow prior:
$x-\tilde{x}^{t}$ $=\left[v_{\phi}(x_{t},t,y)-v_{\phi}(x_{s},s,\emptyset)\right]ds+\eta_{t}$ (17)
[…]
$+\left[v_{\phi}(x_{2\delta_{T}},2\delta_{T},y)-v_{\phi}(x_{\delta_{T}},\delta_{T},\emptyset)\right]dt$
Same as ISM, we ignore the $\eta_{t}$ and use $v_{\phi}(x_{t},t,y)-v_{\phi}(x_{s},s,\emptyset)$ as ISM with Rectified flow prior.
.5 Discussion of the steps of NFE and sample methods
Figure 14: The results of different NFEs and sample methods.
As the number of iterations for push-backward increases, more training time are required. As NFE increases, the LEGO car exhibits more complex structures overall. When viewed vertically, each column represents results of different sampling methods with the same NFE steps. In terms of results for each column, the Euler sampling method demonstrates strong competitiveness, regardless of number of steps.
.6 More Results of FlowDreamer
Figure 15: More results under 3D GS generation setting. Please zoom in for details
Figure 16: More results under NeRF generation setting. Please zoom in for details
As shown in Figure 15 and Figure 16, our FlowDreamer can generate high-fidelity textures and shapes from pretrained rectified flow models both in 3D GS and NeRF. Our method can produce realistic objects, such as sweaters and wooden bowls, including fantastical ones that are rare in reality, like pumpkins with glowing runes and baby dragons.
Figure 17: More qualitative comparison under 3D GS generation setting.
Figure 18: More qualitative comparison under NeRF generation setting.
Comparison with other methods in text-to-3D generation for NeRF and 3D GS, respectively, shows that our FlowDreamer creates 3D objects that match well with the input text prompts, exhibiting high fidelity and intricate details. The coupe is realistic in both color and shape, and the generated watch dial appears lifelike in 3D GS results (See Figure 17). The burrito’s shape and quality are more aligned with the prompts, and the generated tacos are even more realistic in NeRF results (See Figure 18).